Circular No. 5816

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA AQUILAE 1993 G. S. Rossano, R. J. Rudy, R. C. Puetter, and D. K. Lynch, Space and

Environment Technology Center, The Aerospace Corporation, write: “We have observed this object on June 3.5 UT with the Lick Observatory 3-m Shane telescope and the Aerospace cooled scanning germanium spectrometer. The spectrum (range 750-1350 nm, resolution 300-550) is dominated by neutral oxygen at 844.6 and 1128.7 nm excited by the Lyβ flourescence mechanism. Moderate O I at 777.4 and 1316.4 nm is also present. Moderately strong Paβ through Pa_{12} are detected, as well as the Ca II triplet at 849.8, 854.2, and 866.2 nm. He I (1083.0 nm) is weak or absent. The spectrum possesses numerous moderate and weak permitted lines of neutral nitrogen and neutral carbon. Fe II is seen at 892.7, 999.7, 1017.4, 1086.3, and 1112.6 nm. A monochromatic continuum magnitude of 7.0 was observed at 1250 nm.”

C. E. Woodward, University of Wyoming; G. F. Lawrence, R. D. Gehrz, and M. Creese, University of Minnesota; and J. Hakkila, Mankato State University, report the following infrared magnitudes, obtained using the Mt. Lemmon Observatory 1.52-m telecope (+ InSb photometer) on May 28.4 UT, J = 5.91 ± 0.04, H = 5.57 ± 0.04, K = 5.31 ± 0.03, L′ = 4.03 ± 0.04, M = 3.60 ± 0.04; 31.4, 6.28 ± 0.04, 6.10 ± 0.04, 5.82 ± 0.04, 4.67 ± 0.04, 4.17 ± 0.07; June 1.4, 6.32 ± 0.03, 6.13 ± 0.03, 5.83 ± 0.02, 4.69 ± 0.03, 4.21 ± 0.06; 2.5, 6.41 ± 0.04, 6.18 ± 0.06, 5.93 ± 0.04, 4.73 ± 0.05, 4.16 ± 0.06.

SUPERNOVA 1993J IN NGC 3031 R. Pressberger and H. M. Maitzen, Institute for Astronomy, Vienna, re-

port photoelectric measurements of SN 1993J, obtained at the 1-m telescope of the A. Purgathofer Observatory at Klosterneuburg, in the Strömgren system plus intermediate-band (half-width 130 nm) filters g_{1} (503 nm) and g_{2} (521 nm) relative to the comparison star at α = 9^{h}55^{m}41^{s}.5, δ = +69^{o}00′30′′ (equinox 2000.0): Apr. 27.89 UT, u = +1.4, v = +0.43, b = +0.07, g_{1} = –0.10, g_{2} = –0.08, y = –0.47; May 11.91, +2.6, +1.29, +0.87, +0.59, +0.52, +0.17; 14.91, +2.3, +1.28, +0.87, +0.65, +0.58, +0.25; 17.89, +2.4, +1.30, +0.95, +0.68, +0.60, +0.28. The precision of the u measurements is ∼ 0.1 mag, the remaining filters ∼ 0.02 mag.

1993 June 15 Daniel W. E. Green